1.) Assumimg the population proportion of adults who visited a casino in the past 12 months is 0.24, find the approximate probability that the sample proportion of a Gallup poll of 1027 adults who said they visited a casino in the past 12 months is less than 0.22.
Answer #1

Solution

Given that,

=  [p ( 1 - p ) / n] =   [(0.24 * 0.76) / 1027 ] = 0.0133

P( < 0.22) =

= P[( - ) / < (0.22 - 0.24) / 0.0133]

= P(z < -1.50)

= 0.0668

Probability = 0.0668